# The Z-Gromov-Wasserstein Distance

### v1

@@ 3.1.4 The Gromov-Wasserstein distance between Dynamic Metric Spaces
Let $C(\mathbb{R},\mathbb{R}_{\geq 0})$ denote the collection of all continuous functions $f:\mathbb{R}\to\mathbb{R}_{\geq 0}$. For $\lambda\geq 0$ ones defines (see [35, Definition 2.7.1] the $\lambda$-slack interleaving distance between real-valued functions as follows. For $f_{1},f_{2}\in C(\mathbb{R},\mathbb{R}_{\geq 0})$,
@@
According to [35, Defintion 2.9.5], the $(p,\lambda)$-Gromov-Wasserstein distance between two wDMSs $\gamma_{X}$ and $\gamma_{Y}$ is defined as
@@ 3.1.5 The Cut Distance between Graphons
Now, given two probability spaces $(\Omega,\mu)$ and $(\Omega^{\prime},\mu^{\prime})$, and kernels $U$ on $\Omega$ and $U^{\prime}$ on $\Omega^{\prime}$ one defines the cut distance between $(\Omega,U,\mu)$ and $(\Omega^{\prime},U^{\prime},\mu)$ as [32, Theorem 6.9]
@@
While not exactly fitting into our framework, owing to the presence of the cut norm instead of the $L^{p}$ norm, the structure of the cut distance defined above is very closely related to that of the $Z$-Gromov-Wasserstein distance when $Z=\mathbb{R}$. It seems interesting to explore the possibility of formulating an even more general setting than the one in Definition 11 which could (precisely) encompass the cut distance.
@@ 3.2.1 An Alternative Approach to Edge-Attributed Graphs
In Definition 18, we gave a flexible model for graphs with node and edge attributions, which we referred to as attributed networks, following [78]. However, the formalism developed there may be unsatisfactory in practice, as we now explain. Consider a directed graph $(X,E)$, where $X$ is a set of nodes and $E\subset X\times X$ is a set of directed edges (this also captures the notion of an undirected graph, which we consider as a directed graph with a symmetric edge set). Realistic graphical data frequently comes with edge attributes (see the discussion at the beginning of Section 3.1.2)—in practice, this is given by a function from $E$ into some attribute space $(\Omega,d_{\Omega})$. Now observe that the formalism described in Definition 18 (which originates in [78]) models edge attributes as functions of the form $\omega_{X}:X\times X\to\Omega$; that is, the function is not only defined on the edge set, but on the full product space $X\times X$. This difference can be handled in an ad hoc manner: for example, if $\Omega=\mathbb{R}^{d}$, one could assign pairs $(x,x^{\prime})\in(X\times X)\setminus E$ to the zero vector. Such a choice is essentially arbitrary, and does not naturally extend to a more general metric space $\Omega$. We now describe a more principled approach to handling this issue via cone metrics.
Let $(\Omega,d_{\Omega})$ be a metric space, which we will later consider as an edge attribute space. The cone space [11, Section 3.6.2] of $\Omega$ is
@@
where $p_{0}\in\Omega$ is a fixed but arbitrary point. The idea is that this extends the original attribute function $\hat{\omega}_{X}$ (whose domain is $E$) to the full product space $X\times X$, but the cone construction results in the choice of assignment for $(x,x^{\prime})\notin E$ being inconsequential. Node attributes for the graph can be handled as in Definition 18, so that the structure defined here gives a complete refinement of the previous attributed network structure.
@@ 4.1.1 Existence of Optimal Couplings
Our proof strategy is as follows: since $\mathcal{C}(\mu_{X},\mu_{Y})$ is sequentially compact with respect to the weak convergence of probability measures [74, p. 32], it is enough to prove that the distortion functional $\mathrm{dis}_{p}^{Z}$ is continuous in this topology. This approach was also considered in previous works such as [19] and [53], but the same results do not apply in our general setting. For example, [53] heavily utilizes the fact that the kernels are metrics, and [19] uses the fact the bounded continuous $\mathbb{R}$-valued functions are dense in $L^{p}$ spaces, which is not true for a general target metric space (consider $Z=\{0,1\}$). The corresponding result for the fused network GW distances (Definition 20) was established in [78]. In light of Proposition 23, and the fact that [78] considered more restrictive structures with bounded and continuous kernels, our result further generalizes the result of [78, Theorem 2.4].
To deal with our case, we utilize the following lemma, which generalizes [59, Lemma 1.8] to the Gromov-Wasserstein setting.
Lemma 27.
@@
As mentioned earlier, it is enough to prove that $\mathcal{C}(\mu_{X},\mu_{Y})\ni\pi\mapsto\mathrm{dis}_{p}^{Z}(\pi)$ is continuous with respect to the topology of weak convergence. We can apply Lemma 27 by setting
@@
so the assumption $c(a,b)\leq f(a)+g(b)$ is satisfied. The integrability of $f\circ a$ and $g\circ b$ is satisfied by the $L^{p}$ assumption on $a=\omega_{X}$ and $b=\omega_{Y}$, so all assumptions are clear. Thus, $\mathrm{dis}_{p}$ is continuous, and optimal couplings exist for $1\leq p<\infty$. Since the $p=\infty$ case can be seen as the supremum of $\mathrm{dis}_{p}$ for $p<\infty$, we have lower semicontinuity, and the optimal coupling exists for $p=\infty$. ∎
@@ 4.1.2 $Z$-Gromov-Wasserstein Distance is a Metric
Definition 28.
We define a pair of $(Z,p)$-networks $X=(X,\omega_{X},\mu_{X})$ and $(Y,\omega_{Y},\mu_{Y})$ to be weakly isomorphic if there exists a $(Z,p)$-network $W=(W,\omega_{W},\mu_{W})$, and maps $\phi_{X}:W\to X$ and $\phi_{Y}:W\to Y$ such that
@@
For the Fused GW [72] and Fused Network GW [78], it was previously only shown that they satisfy “relaxed” triangle inequalities, including an extra scaling term depending on the $q$ parameter. With our slight reformulation of these distances (see Remark 21 and Remark 22), we have the following corollary of Theorem 29. The result follows immediately by transporting the definition of weak isomorphism into the appropriate context and via Proposition 23 and Theorem 29.
Corollary 30.
@@
Lemma 31 (Gluing Lemma, [74]).
@@
First, it is clear that $\mathrm{GW}_{p}^{Z}(X,Y)\geq 0$ for any $Z$-networks $X,Y$. For symmetry, the symmetry of $d$ proves that the mapping $f:X\times Y\to Y\times X,f(x,y)=(y,x)$ defines a bijection $\mathcal{C}(\mu_{X},\mu_{Y})\ni\pi\mapsto f_{*}\pi\in\mathcal{C}(\mu_{Y},\mu_{X})$ such that $\mathrm{dis}_{p}^{Z}(\mu)=\mathrm{dis}_{p}^{Z}(f_{*}\pi)$, and taking the infimum shows that $\mathrm{GW}_{p}^{Z}(X,Y)=\mathrm{GW}_{p}^{Z}(Y,X)$.
@@ 4.1.3 $Z$-Gromov-Wasserstein Space as a Quotient of the $L^{p}$ Space
Definition 33 (Invariant Transforms, [67]).
We say that a Borel measurable map $\phi:[0,1]\to[0,1]$ is $\mathscr{L}$-invariant if $\phi_{*}\mathscr{L}=\mathscr{L}$. Define the semigroup of all $\mathscr{L}$-invariant maps by function composition by $\textnormal{Inv}([0,1],\mathscr{L})$.
@@
where $D_{p}$ is the $L^{p}$ distance (see Definition 6) and this family of metric spaces works as a representation of the GW spaces:
@@
As the proof in [67, Theorem 5.10] does not rely on the properties of the target metric space, trivial modifications allow us to apply the proof to our case. ∎
@@ 4.2.1 Separability
Here, $D_{p}$ denotes the $L^{p}$ distance of Definition 6.
@@
Proposition 36.
@@
Without loss of generality, we consider $Z$-networks of the form $([0,1],\omega,\mathscr{L})$ where $\mathscr{L}$ is the Lebesgue measure by Proposition 32. Since $L^{p}([0,1]\times[0,1],\mathscr{L}\otimes\mathscr{L};Z)$ is separable by Proposition 8, we can take a countable dense subset $\{\omega_{i}\}_{i=1}^{\infty}$ of $L^{p}([0,1]\times[0,1],\mathscr{L}\otimes\mathscr{L};Z)$. Now, for any $Z$-network $X=([0,1],\omega,\mathscr{L})$ and any $\epsilon>0$, there exists $\omega_{i}$ such that $D_{p}(\omega,\omega_{i})<\epsilon$. By constructing a $Z$-network by $X_{i}=([0,1],\omega_{i},\mathscr{L})$, and by Lemma 35, we see that $\mathrm{GW}_{p}^{Z}(X,X_{i})<\epsilon$. Therefore, $\{X_{i}\}_{i=1}^{\infty}$ is dense in $\mathcal{M}_{\sim}^{Z,p}$. ∎
@@
We first take an arbitrary $(Z,p)$-network. By parameterization (Proposition 32), it is weakly isomorphic to a $(Z,p)$-network of the form $X=([0,1],\omega,\mathscr{L})$. The kernel $\omega$ belongs to $L^{p}([0,1]^{2},\mathscr{L}^{2};Z)$, so by Proposition 8, it can be approximated by a piecewise constant function $\omega_{n}$ that is constant on a grid of step size $1/n$ with $n$ small enough. That is, we can partition $[0,1]$ into $n$ pieces $R_{i}=[i/n,(i+1)/n)$ for $i=0,\ldots,n-2$, $R_{n-1}=[(n-1)/n,n]$ and take $\omega_{n}$ to be constant on each $R_{i}\times R_{j}$. The $Z$-network $X_{n}=([0,1],\omega_{n},\mathscr{L})$ is weakly isomorphic to ($(\{1,\cdots,n\},\Omega_{n},\frac{1}{n}\sum_{i=1}^{n}\delta_{i})$ where $\Omega_{n}$ is the $n\times n$, $Z$-valued matrix with $\Omega_{n}(i,j)=\omega_{n}(i/n,j/n)$ because we can construct a measure preserving map $\phi:X\to\{1,\cdots,n\}$ by $\phi(x)=i$ whenever $x\in R_{i}$, and this map satisfies $\phi^{*}\Omega_{n}=\omega_{n}$ by definition. Finally, by using Lemma 35 to $X$ and $X_{n}$, we have the density. ∎
From this, we can provide an alternative proof of the separability of $\mathcal{M}_{\sim}^{Z,p}$ since each of $\mathcal{M}_{n}^{Z}$ is separable under $\mathrm{GW}_{p}^{Z}$. This is because this space consists of $(Z,p)$-networks of the form ($(\{1,\cdots,n\},\omega,\frac{1}{n}\sum_{i=1}^{n}\delta_{i})$ i.e. everything except the kernel is the same, so the proof of Proposition 36 works the same by replacing $[0,1]$ to $\{1,\cdots,n\}$ and $\mathscr{L}$ to $\frac{1}{n}\sum_{i=1}^{n}\delta_{i}$.
@@ 4.2.2 Completeness
Suppose $Z$ is complete. Then, the proof that $\mathrm{GW}_{p}^{Z}$ is complete follows exactly as in the case $Z=\mathbb{R}$ [67, Theorem 5.8], since $L^{p}(X\times X,\mu_{X}\otimes\mu_{X};Z)$ is complete (Proposition 8). This completes one direction of the theorem.
Let us now consider the converse. We will generalize the proof for the Wasserstein space [27] to our general setting. Suppose that $\mathrm{GW}_{p}^{Z}$ is complete. Let $\{z_{n}\}$ be a Cauchy sequence in $Z$. We will show that $\{z_{n}\}$ converges to some $z\in Z$. For each $n$, let $X_{n}=(\{0\},\omega_{n},\delta_{0})$ where $\omega_{n}:\{0\}\times\{0\}\to Z$ is the function $\omega_{n}(0,0)=z_{n}$. Then, $\{X_{n}\}$ is a Cauchy sequence in $\mathcal{M}^{Z}_{\sim}$ with respect to $\mathrm{GW}_{p}^{Z}$ because $\mathrm{GW}_{p}^{Z}(X_{n},X_{m})=\frac{1}{2}d_{Z}(z_{n},z_{m})$. By completeness, there exists $X=(X,\omega,\mu)\in\mathcal{M}^{Z}_{\sim}$ such that $\mathrm{GW}_{p}^{Z}(X_{n},X)\to 0$. Since the only coupling with a Dirac measure is the product measure, we can explicitly calculate $\mathrm{GW}_{p}^{Z}(X_{n},X)$ as
@@
For brevity, we denote $\omega_{*}(\mu\otimes\mu)$ by $\mu_{\omega}$. We note that $\mu_{\omega}$ is a probability measure on $Z$ since $\mu\otimes\mu$ is a probability measure on $X\times X$. Therefore, Markov’s inequality is applicable to show that
@@
which shows that $\mu_{\omega}(B_{\epsilon}(z_{n}))\to 1$ as $n\to\infty$ for any $\epsilon>0$. Here, $B_{\epsilon}(z_{n})=\{z\in Z:d_{Z}(z_{n},z)<\epsilon\}$. Intuitively, the above means that the mass of $\mu_{\omega}$ is concentrated around $z_{n}$ as $n\to\infty$, or in other words, $z_{n}$ gets closer and closer to the support point of $\mu_{\omega}$. Indeed, for any $z\in\textrm{supp}(\mu_{\omega})$, we have $\mu_{\omega}(B_{\epsilon}(z))>0$ for any $\epsilon>0$ by definition. This means that, if $n$ is large enough, $B_{\epsilon}(z)$ and $B_{\epsilon}(z_{n})$ should have a nonempty intersection for any $\epsilon>0$ because $\mu_{\omega}(B_{\epsilon}(z))>0$ is a constant and $\mu_{\omega}(B_{\epsilon}(z_{n}))\to 1$, so the sum of two measures will be larger than $1$ at some point. Thus, for any $\epsilon>0$, there exists a natural number $N$ such that if $n\geq N$, we can take $v\in\textrm{supp}(\mu_{\omega})\cap B_{\epsilon}(z_{n})$ so that $d_{Z}(z,z_{n})\leq d_{Z}(z,v)+d_{Z}(v,z_{n})<2\epsilon$. Since $\epsilon>0$ is arbitrary, This shows that $z_{n}\to z$ as $n\to\infty$. As any Cauchy sequence converges, we have shown that $Z$ is complete. ∎

### v2

@@ 3.1.4 The Gromov-Wasserstein distance between Dynamic Metric Spaces
Let $C(\mathbb{R},\mathbb{R}_{\geq 0})$ denote the collection of all continuous functions $f:\mathbb{R}\to\mathbb{R}_{\geq 0}$. For $\lambda\geq 0$ ones defines (see [43, Definition 2.7.1] the $\lambda$-slack interleaving distance between real-valued functions as follows. For $f_{1},f_{2}\in C(\mathbb{R},\mathbb{R}_{\geq 0})$,
@@
According to [43, Defintion 2.9.5], the $(p,\lambda)$-Gromov-Wasserstein distance between two wDMSs $\gamma_{X}$ and $\gamma_{Y}$ is defined as
@@ 3.1.5 The Cut Distance between Graphons
Now, given two probability spaces $(\Omega,\mu)$ and $(\Omega^{\prime},\mu^{\prime})$, and kernels $U$ on $\Omega$ and $U^{\prime}$ on $\Omega^{\prime}$ one defines the cut distance between $(\Omega,U,\mu)$ and $(\Omega^{\prime},U^{\prime},\mu)$ as [40, Theorem 6.9]
@@
While not exactly fitting into our framework, owing to the presence of the cut norm instead of the $L^{p}$ norm, the structure of the cut distance defined above is very closely related to that of the $Z$-Gromov-Wasserstein distance when $Z=\mathbb{R}$. It seems interesting to explore the possibility of formulating an even more general setting than the one in Definition 2.11 which could (precisely) encompass the cut distance.
@@ 3.2.1 An Alternative Approach to Edge-Attributed Graphs
In Definition 3.6, we gave a flexible model for graphs with node and edge attributions, which we referred to as attributed networks, following [90]. However, the formalism developed there may be unsatisfactory in practice, as we now explain. Consider a directed graph $(X,E)$, where $X$ is a set of nodes and $E\subset X\times X$ is a set of directed edges (this also captures the notion of an undirected graph, which we consider as a directed graph with a symmetric edge set). Realistic graphical data frequently comes with edge attributes (see the discussion at the beginning of Section 3.1.2)—in practice, this is given by a function from $E$ into some attribute space $(\Omega,d_{\Omega})$. Now observe that the formalism described in Definition 3.6 (which originates in [90]) models edge attributes as functions of the form $\omega_{X}:X\times X\to\Omega$; that is, the function is not only defined on the edge set, but on the full product space $X\times X$. This difference can be handled in an ad hoc manner: for example, if $\Omega=\mathbb{R}^{d}$, one could assign pairs $(x,x^{\prime})\in(X\times X)\setminus E$ to the zero vector. Such a choice is essentially arbitrary, and does not naturally extend to a more general metric space $\Omega$. We now describe a more principled approach to handling this issue via cone metrics.
Let $(\Omega,d_{\Omega})$ be a metric space, which we will later consider as an edge attribute space. The cone space [13, Section 3.6.2] of $\Omega$ is
@@
where $p_{0}\in\Omega$ is a fixed but arbitrary point. The idea is that this extends the original attribute function $\hat{\omega}_{X}$ (whose domain is $E$) to the full product space $X\times X$, but the cone construction results in the choice of assignment for $(x,x^{\prime})\notin E$ being inconsequential. Node attributes for the graph can be handled as in Definition 3.6, so that the structure defined here gives a complete refinement of the previous attributed network structure.
@@ 4.1.1 Existence of Optimal Couplings
Our proof strategy is as follows: since $\mathcal{C}(\mu_{X},\mu_{Y})$ is sequentially compact with respect to the weak convergence of probability measures [86, p. 32], it is enough to prove that the distortion functional $\mathrm{dis}_{p}^{Z}$ is continuous in this topology. This approach was also considered in previous works such as [23] and [62], but the same results do not apply in our general setting. For example, [62] heavily utilizes the fact that the kernels are metrics, and [23] uses the fact the bounded continuous $\mathbb{R}$-valued functions are dense in $L^{p}$ spaces, which is not true for a general target metric space (consider $Z=\{0,1\}$). The corresponding result for the fused network GW distances (Definition 3.8) was established in [90]. In light of Proposition 3.11, and the fact that [90] considered more restrictive structures with bounded and continuous kernels, our result further generalizes the result of [90, Theorem 2.4].
To deal with our case, we utilize the following lemma, which generalizes [69, Lemma 1.8] to the Gromov-Wasserstein setting.
Lemma 4.1.
@@
As mentioned earlier, it is enough to prove that $\mathcal{C}(\mu_{X},\mu_{Y})\ni\pi\mapsto\mathrm{dis}_{p}^{Z}(\pi)$ is continuous with respect to the topology of weak convergence. We can apply Lemma 4.1 by setting
@@
so the assumption $c(a,b)\leq f(a)+g(b)$ is satisfied. The integrability of $f\circ a$ and $g\circ b$ is satisfied by the $L^{p}$ assumption on $a=\omega_{X}$ and $b=\omega_{Y}$, so all assumptions are clear. Thus, $\mathrm{dis}_{p}$ is continuous, and an optimal coupling exists for $1\leq p<\infty$. Since the $p=\infty$ case can be seen as the supremum of $\mathrm{dis}_{p}$ for $p<\infty$, we have lower semicontinuity, and an optimal coupling exists for $p=\infty$. ∎
@@ 4.1.2 $Z$-Gromov-Wasserstein Distance is a Metric
Definition 4.2.
We define a pair of $(Z,p)$-networks $X=(X,\omega_{X},\mu_{X})$ and $Y=(Y,\omega_{Y},\mu_{Y})$ to be weakly isomorphic if there exists a $(Z,p)$-network $W=(W,\omega_{W},\mu_{W})$, and maps $\phi_{X}:W\to X$ and $\phi_{Y}:W\to Y$ such that
@@
For the Fused GW [84] and Fused Network GW [90], it was previously only shown that they satisfy “relaxed” triangle inequalities, including an extra scaling term depending on the $q$ parameter. With our slight reformulation of these distances (see Remark 3.9 and Remark 3.10), we have the following corollary of Theorem 3. The result follows immediately by transporting the definition of weak isomorphism into the appropriate context and via Proposition 3.11 and Theorem 3.
Corollary 4.3.
@@
Lemma 4.4 (Gluing Lemma, [86]).
@@
First, it is clear that $\mathrm{GW}_{p}^{Z}(X,Y)\geq 0$ for any $Z$-networks $X,Y$. For symmetry, the symmetry of $d$ proves that the mapping $f:X\times Y\to Y\times X,f(x,y)=(y,x)$ defines a bijection $\mathcal{C}(\mu_{X},\mu_{Y})\ni\pi\mapsto f_{*}\pi\in\mathcal{C}(\mu_{Y},\mu_{X})$ such that $\mathrm{dis}_{p}^{Z}(\pi)=\mathrm{dis}_{p}^{Z}(f_{*}\pi)$, and taking the infimum shows that $\mathrm{GW}_{p}^{Z}(X,Y)=\mathrm{GW}_{p}^{Z}(Y,X)$.
@@ 4.1.3 $Z$-Gromov-Wasserstein Space as a Quotient of the $L^{p}$ Space
Definition 4.6 (Invariant Transforms, [77]).
We say that a Borel measurable map $\phi:[0,1]\to[0,1]$ is $\mathscr{L}$-invariant if $\phi_{*}\mathscr{L}=\mathscr{L}$. The $\mathscr{L}$-invariant maps form a semigroup via function composition, which we denote by $\textnormal{Inv}([0,1],\mathscr{L})$.
@@
where $D_{p}$ is the $L^{p}$ distance (see Definition 2.6) and this family of metric spaces works as a representation of the GW spaces:
@@
As the proof in [77, Theorem 5.10] does not rely on the properties of the target metric space, trivial modifications allow us to apply the proof to our case. ∎
@@ 4.2.1 Separability
Here, $D_{p}$ denotes the $L^{p}$ distance of Definition 2.6.
@@
for $p\in[1,\infty)$. The $p=\infty$ case is handled by replacing the integral with an essential supremum. ∎
Proposition 4.8.
@@
Without loss of generality, we consider $Z$-networks of the form $([0,1],\omega,\mathscr{L})$ where $\mathscr{L}$ is the Lebesgue measure by Proposition 4.5. Since $L^{p}([0,1]\times[0,1],\mathscr{L}\otimes\mathscr{L};Z)$ is separable by Proposition 2.8, we can take a countable dense subset $\{\omega_{i}\}_{i=1}^{\infty}$ of $L^{p}([0,1]\times[0,1],\mathscr{L}\otimes\mathscr{L};Z)$. Now, for any $Z$-network $X=([0,1],\omega,\mathscr{L})$ and any $\epsilon>0$, there exists $\omega_{i}$ such that $D_{p}(\omega,\omega_{i})<\epsilon$. By constructing a $Z$-network by $X_{i}=([0,1],\omega_{i},\mathscr{L})$, and by Lemma 4.7, we see that $\mathrm{GW}_{p}^{Z}(X,X_{i})<\epsilon$. Therefore, $\{X_{i}\}_{i=1}^{\infty}$ is dense in $\mathcal{M}_{\sim}^{Z,p}$. ∎
@@
We first take an arbitrary $(Z,p)$-network. By parameterization (Proposition 4.5), it is weakly isomorphic to a $(Z,p)$-network of the form $X=([0,1],\omega,\mathscr{L})$. The kernel $\omega$ belongs to $L^{p}([0,1]^{2},\mathscr{L}^{2};Z)$, so by Proposition 2.8, it can be approximated by a piecewise constant function $\omega_{n}$ that is constant on a grid of step size $1/n$ with $n$ small enough. That is, we can partition $[0,1]$ into $n$ pieces $R_{i}=[i/n,(i+1)/n)$ for $i=0,\ldots,n-2$, $R_{n-1}=[(n-1)/n,n]$ and take $\omega_{n}$ to be constant on each $R_{i}\times R_{j}$. The $Z$-network $X_{n}=([0,1],\omega_{n},\mathscr{L})$ is weakly isomorphic to ($(\{1,\cdots,n\},\Omega_{n},\frac{1}{n}\sum_{i=1}^{n}\delta_{i})$ where $\Omega_{n}$ is the $n\times n$, $Z$-valued matrix with $\Omega_{n}(i,j)=\omega_{n}(i/n,j/n)$ because we can construct a measure preserving map $\phi:X\to\{1,\cdots,n\}$ by $\phi(x)=i$ whenever $x\in R_{i}$, and this map satisfies $\phi^{*}\Omega_{n}=\omega_{n}$ by definition. Finally, by using Lemma 4.7 to $X$ and $X_{n}$, we have the density. ∎
From this, we can provide an alternative proof of the separability of $\mathcal{M}_{\sim}^{Z,p}$ since each of $\mathcal{M}_{n}^{Z}$ is separable under $\mathrm{GW}_{p}^{Z}$. This is because this space consists of $(Z,p)$-networks of the form ($(\{1,\cdots,n\},\omega,\frac{1}{n}\sum_{i=1}^{n}\delta_{i})$ i.e. everything except the kernel is the same, so the proof of Proposition 4.8 works the same by replacing $[0,1]$ to $\{1,\cdots,n\}$ and $\mathscr{L}$ to $\frac{1}{n}\sum_{i=1}^{n}\delta_{i}$.
@@ 4.2.2 Completeness
Suppose $Z$ is complete. Then, the proof that $\mathrm{GW}_{p}^{Z}$ is complete follows exactly as in the case $Z=\mathbb{R}$ [77, Theorem 5.8], since $L^{p}(X\times X,\mu_{X}\otimes\mu_{X};Z)$ is complete (Proposition 2.8). This completes one direction of the theorem. Let us now consider the converse. We will generalize the proof for the Wasserstein space [35] to our setting. Suppose that $\mathrm{GW}_{p}^{Z}$ is complete. Let $\{z_{n}\}$ be a Cauchy sequence in $Z$. We will show that $\{z_{n}\}$ converges to some $z\in Z$. For each $n$, let $X_{n}=(\{0\},\omega_{n},\delta_{0})$ where $\omega_{n}:\{0\}\times\{0\}\to Z$ is the function $\omega_{n}(0,0)=z_{n}$. Then, $\{X_{n}\}$ is a Cauchy sequence in $\mathcal{M}^{Z}_{\sim}$ with respect to $\mathrm{GW}_{p}^{Z}$ because $\mathrm{GW}_{p}^{Z}(X_{n},X_{m})=\frac{1}{2}d_{Z}(z_{n},z_{m})$. By completeness, there exists $X=(X,\omega,\mu)\in\mathcal{M}^{Z}_{\sim}$ such that $\mathrm{GW}_{p}^{Z}(X_{n},X)\to 0$. Since the only coupling with a Dirac measure is the product measure, we can explicitly calculate $\mathrm{GW}_{p}^{Z}(X_{n},X)$ as
@@
for $p<\infty$. For $p=\infty$, we have
$\mathrm{GW}_{\infty}^{Z}(X_{n},X)=\frac{1}{2}\|d(z_{n},z)\|_{L^{\infty}(\omega_{*}(\mu\otimes\mu))}$ (67)
For brevity, we denote $\omega_{*}(\mu\otimes\mu)$ by $\mu_{\omega}$. We note that $\mu_{\omega}$ is a probability measure on $Z$ since $\mu\otimes\mu$ is a probability measure on $X\times X$. Therefore, if $p<\infty$, Markov’s inequality is applicable to show that
@@
Moreover, if $p=\infty$, by the monotonicity of $L^{p}$ norms,
$\epsilon(1-\mu_{\omega}(B_{\epsilon}(z_{n}))=\epsilon\mu_{\omega}(\{z\in Z:d_{Z}(z_{n},z)\geq\epsilon\})\leq\|d_{Z}(z_{n},z)\|_{L^{1}(\mu_{\omega})}\leq\|d_{Z}(z_{n},z)\|_{L^{\infty}(\mu_{\omega})}$ (69)
which shows that $\mu_{\omega}(B_{\epsilon}(z_{n}))\to 1$ as $n\to\infty$ for any $\epsilon>0$. Here, $B_{\epsilon}(z_{n})=\{z\in Z:d_{Z}(z_{n},z)<\epsilon\}$.
Intuitively, the above means that the mass of $\mu_{\omega}$ is concentrated around $z_{n}$ as $n\to\infty$, or in other words, $z_{n}$ gets closer and closer to the support point of $\mu_{\omega}$. Indeed, for any $z\in\textrm{supp}(\mu_{\omega})$, we have $\mu_{\omega}(B_{\epsilon}(z))>0$ for any $\epsilon>0$ by definition. This means that, if $n$ is large enough, $B_{\epsilon}(z)$ and $B_{\epsilon}(z_{n})$ should have a nonempty intersection for any $\epsilon>0$ because $\mu_{\omega}(B_{\epsilon}(z))>0$ is a constant and $\mu_{\omega}(B_{\epsilon}(z_{n}))\to 1$, so the sum of two measures will be larger than $1$ at some point. Thus, for any $\epsilon>0$, there exists a natural number $N$ such that if $n\geq N$, we can take $v\in\textrm{supp}(\mu_{\omega})\cap B_{\epsilon}(z_{n})$ so that $d_{Z}(z,z_{n})\leq d_{Z}(z,v)+d_{Z}(v,z_{n})<2\epsilon$. Since $\epsilon>0$ is arbitrary, This shows that $z_{n}\to z$ as $n\to\infty$. As any Cauchy sequence converges, we have shown that $Z$ is complete. ∎